OpenStudy (anonymous):

Solve the equation |3x-6|=15

Parth (parthkohli):

\( \color{Black}{\Rightarrow 3x - 6 = \pm 15 }\) Can you do it now?

Parth (parthkohli):

Basically the distance of 3x - 6 from 0 is 15. Both 15 and -15 are 15 units away.

Parth (parthkohli):

We have two equations here. \( \color{Black}{\Rightarrow 3x - 6 = 15 }\) \( \color{Black}{\Rightarrow 3x - 6 = -15 }\) Solve both to get the two solutions